OpenStudy (anonymous):

Write equivalent fractions for 3/5, 3/4, and 2/3 using the least common denominator.

myininaya (myininaya):

5*4*3=60 is lcm \[\frac{3}{5} \cdot \frac{12}{12}=\frac{36}{60}, \frac{3}{4} \cdot \frac{15}{15}=\frac{45}{60},\frac{2}{3} \cdot \frac{20}{20}=\frac{40}{60}\]

jimthompson5910 (jim_thompson5910):

LCM of 5 and 4 is 20. The LCM of 20 and 3 is 60 So the LCM of 5, 4, and 3 is 60 So the LCD is 60. This means that we must get all denominators equal to 60. We do this by multiplying top and bottom of 3/5 by 12 to get 36/60 and by multiplying top and bottom of 3/4 by 15 to get 45/60 and by multiplying top and bottom of 2/3 by 20 to get 40/60 So the three fractions turn into: 36/60, 45/60, and 40/60

OpenStudy (anonymous):

Thanksss :D